# Gradient-flow SDEs have unique transient population dynamics

### v2

@@ A.1 The distributional definition of the Fokker-Planck equation
In this setting, the existence and uniqueness of a weak solution follows from [Bogachev et al., 2022, Theorem 9.4.8, Example 9.4.7], which applies because $\sigma^{2}$ is constant and we assumed that $\nabla\Psi$ is Lipschitz and obeys the growth condition, $\|\nabla\Psi(x)\|\leq K(1+\|x\|)$. This result is also given in Stroock [2008, Theorem 1.1.9], which additionally shows that the family $(\mu_{t})_{t\in(0,T)}$ is continuous, under the topology of weak convergence. Thus, the evolution of marginals can be identified with a continuous transition probability function, even under the weak formulation.
@@ A.2 Additional proofs
The following results are used to complete the proof of ˜4.3. Since the observed marginals in ˜4.3 are sampled after some time $T_{1}>0$, we note that it suffices to consider times $t\geq T_{1}>0$. Also recall that for this result, we assume that $\Psi\in C^{\infty}(\mathbb{R}^{d})$. This facilitates the proof, by ensuring that any marginal $p(\cdot,t)$ of a candidate gradient-flow SDE exhibits nice regularity and decay. Indeed, by elliptic regularity theory, the Fokker-Planck operator $\mathcal{L}_{\shortminus\nabla\Psi,\sigma^{2}}$ smooths marginals within any finite time.
@@
As defined in ˜4.3, let $\mathcal{I}_{i}=\{t\in[T_{i},T_{i+1}]\mid p(\cdot,t)=q(\cdot,t)\}$, and let $\{t_{n}\}_{n\geq 0}$ be a sequence of times in $\mathcal{I}_{i}$, which converges to $t_{i}^{*}$. Then, $p(\cdot,t_{i}^{*})=q(\cdot,t_{i}^{*})$ and $\frac{\partial}{\partial_{t}}p(\cdot,t_{i}^{*})=\frac{\partial}{\partial_{t}}q(\cdot,t_{i}^{*})$.
@@
Let $(p(\cdot,t))_{t\geq T_{1}}$ be the marginals of a Langevin SDE (1) from the earliest possible observation time in the setting of ˜4.3, and suppose that $\Psi\in C^{\infty}(\mathbb{R}^{d})$. Then, the free energy $J(\rho)=\int_{\mathbb{R}^{d}}\Psi(x)\rho(x)\mathrm{d}x+\beta\int_{\mathbb{R}^{d}}\rho(x)\log\rho(x)\mathrm{d}x$, such that $\beta=\frac{\sigma^{2}}{2}$, strictly decreases in time, unless the process is at its stationary Gibbs distribution, $p_{\mathrm{eq}}$. Thus, $p_{\mathrm{eq}}$ is the only possible marginal that can repeat at distinct times.
@@
$\frac{d}{dt}J(p(\cdot,t))$ $=\int_{\mathbb{R}^{d}}\bigl(\Psi(x)+\beta(\log p(x,t)+1)\bigr)\,\frac{\partial}{\partial_{t}}p(x,t)\,\mathrm{d}x$ (22)
@@ Appendix B Identifiability in the time-inhomogeneous case
We apply Theorem˜4.2 on each regime $t_{i},t_{i+1})$. Since the regimes form a partition of $[0,T]$, this guarantees that there is a unique SDE with the observed marginals $\bigl(p(\cdot,t\bigr)_{t\in[0,T]}$. ∎
@@ Appendix C Additional non-identifiability examples
For gradient-flow SDEs, we have proven that non-identifiability from marginals arises if and only the observed marginals also correspond to the marginals of a residual process in equilibrium. An example for the time-homogeneous case was given in ˜3.2, such that the equilibrium marginals are constant, and an example for the time-inhomogenous case was given in ˜B.3, such that the equilibrium distribution of the residual process continuously changes in time.
@@
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}-10&0\\ 0&b\end{bmatrix}\mathrm{d}t+\begin{bmatrix}\sqrt{10}&0\\ 0&c\end{bmatrix}\mathrm{d}W_{t},$ $Y_{0}\sim\mathcal{N}(0,1)\times[a,b]$ (30)
@@
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}Y_{t}\>\mathrm{d}t+\mathrm{d}W_{t},$ $Y_{0}\sim\mathcal{N}(0,I)$
@@
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}1/3&4/3\\ 2/3&-1/3\end{bmatrix}Y_{t}\>\mathrm{d}t+\begin{bmatrix}1&2\\ -1&-2\end{bmatrix}\mathrm{d}W_{t},$ $Y_{0}=\begin{bmatrix}1\\ -1\end{bmatrix}$
Appendix D Convergence and runtime of nn-APPEX
@@
By Girsanov’s theorem, (31) is only finite for $Q$ and $P$, if their underlying SDEs share the same diffusion [Vargas et al., 2021]. Besides non-identifiability, this is another cited reason for why previous Schrodinger Bridge methods assume that diffusivity must be fixed [Shen et al., 2025]. However, in the practical setting, we only observe a finite number of marginals, so we should instead optimize $\mathrm{KL}$ divergence over the discretized path space $\Omega_{N}=C(\{t_{i}\}_{i=1}^{N},\mathbb{R}^{d})$. Then, for each law on paths, we only observe the couplings, $Q_{t_{i},t_{i+1}}$ and $P_{t_{i},t_{i+1}}$, between consecutive times. Let $Q^{N}$ and $P^{N}$ denote the concatenations of these couplings from $t_{1}$ to $t_{N}$. Then, by Benamou et al. [2019][Lemma 3.4], evaluating the $\mathrm{KL}$ divergence over $\Omega_{N}=C(\{t_{i}\}_{i=1}^{N},\mathbb{R}^{d})$ yields
@@
As noted in Section˜5, the iterates will produce a monotonically decreasing sequence of $\mathrm{KL}$ divergences between reconstructed laws and MLE estimated laws. With the above argument, we have that the $\mathrm{KL}$ divergences are bounded above by some $M>0$, and from below by 0 (attained by the true gradient-flow SDE parameters $(-\nabla\Psi,\sigma^{2})$). Thus, the sequence of $\mathrm{KL}$ divergences must converge. However, as noted in Shen et al. [2025], this does not imply that the arguments necessarily converge, since multiple pairs $(P,Q)$ may produce the same divergence, such that the iterative algorithm gets stuck in a cycle. While this has not been observed empirically, and is impossible if the divergence is 0 (since there is a unique gradient-flow SDE satisfying the marginal constraints), it is an open question whether the iterative scheme will converge to the unique optimal parameters. Indeed, while convergence of the iterative SB refinement algorithm has been proven in the case where the family of admissible probability transition densities is convex [Shen et al., 2025][Proposition 1], this condition does not generally hold, and is false for the family of gradient-flow SDEs.
D.1 Runtime
@@
We note that nn-APPEX’s runtime will vary for different implementations of these subprocedures. In particular, there are multiple approaches for solving the multi-marginal SB problem. One approach is to consider all contiguous pairs of marginals, and then to apply iterative proportional fitting (Sinkhorn’s algorithm) on each pair to determine the distribution over couplings [Shen et al., 2025, Guan et al., 2024]. To enforce consistency across transitions, we instead apply a multi-marginal iterative proportional fitting across all marginals, see [Marino and Gerolin, 2020][(4.9)]. This algorithm is analogous to the two-marginals setting, but it instead rescales per-time slices for each marginal, rather than just the endpoints. We observed better accuracy for this approach, given the same stopping criteria and maximum number of iterations, and were thus able to reduce runtime with relatively fewer iterations. We similarly note that MLE drift estimation can be implemented using different neural network architectures and hyperparameters. The details of our optimization are provided in Section˜6. In contrast, the diffusion MLE estimate is a closed formula, and is thus cheap to compute. We report the average runtime (in seconds) of each of the three subprocedures for a single iteration, run on our main experiments:
@@ Appendix E Maximum likelihood estimation for drift and diffusion
If we observe $N$ independent trajectories over $T$ time steps ($T+1$ observed times), let $X=\bigl\{x_{k\Delta t}^{(n)},x_{(k+1)\Delta t}^{(n)}:n\in\{1,\cdots,N\},k\in\{0,\cdots,T-1\}\bigr\}$. We similarly observe that the negative log-likelihood function is given by
$l\left(\theta,\sigma^{2}|X\right)=-\log\left(\prod_{n=1}^{N}\prod_{k=0}^{T-1}p(x_{(k+1)\Delta t}^{(n)}|x_{k\Delta t}^{(n)})\right)=-\sum_{n=1}^{N}\sum_{k=0}^{T-1}\log(p(x_{(k+1)\Delta t}^{(n)}|x_{k\Delta t}^{(n)})),$
@@
$\ell(\theta)=\sum_{n=1}^{N}\sum_{k=0}^{T-1}\Big\|x_{(k+1)\Delta t}^{(n)}-x_{k\Delta t}^{(n)}+\nabla\Psi_{\theta}(x_{k\Delta t}^{(n)})\Delta t\Big\|_{2}^{2}.$ (35)
@@
$0$ $=\sum_{n=1}^{N}\sum_{k=0}^{T-1}\frac{d}{2\sigma^{2}}-\frac{1}{2\sigma^{4}\Delta t}\sum_{n=1}^{N}\sum_{k=0}^{T-1}\Big\|x_{(k+1)\Delta t}^{(n)}-x_{k\Delta t}^{(n)}+\nabla\Psi_{\hat{\theta}}(x_{k\Delta t}^{(n)})\Delta t\Big\|_{2}^{2}$

### v3

@@ Acknowledgments.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes/No/Not Applicable]
      The data-generating model is defined in (1) and precise mathematical details and blanket assumptions are given in Section˜2. Our inference algorithm is defined in Section˜4 and further details are given in Appendix˜D
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes/No/Not Applicable]
      The mathematical properties of our algorithm are discussed in Section˜4, with related derivations in Appendix˜D and Appendix˜E. Relevant numerical details are provided in Section˜5, and runtimes are reported in Table˜3.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes/No/Not Applicable]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes/No/Not Applicable]
      All assumptions, outside the blanket assumptions (two textbook assumptions required for SDE existence and uniqueness), are contained within the statements of the theoretical result.
    2. (b)
      Complete proofs of all theoretical results. [Yes/No/Not Applicable]
      Given the importance of our theory and the relative brevity of the proofs, we include full proofs in Section˜3, under Theorem˜3.2 and ˜3.3 respectively. Technical details related to our inference algorithm are derived in Appendix˜D and Appendix˜E.
    3. (c)
      Clear explanations of any assumptions. [Yes/No/Not Applicable]
      We discuss assumptions, their relevance for practical inference, and comparison with assumptions from previous literature throughout the paper. For example, see the discussion after the proof of Theorem˜3.2.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes/No/Not Applicable]
      A zip file of the code repository with scripts for running the experiments, and a detailed README file are included in supplemental material.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes/No/Not Applicable] We report implementation details in Section˜5 and emphasize that we use exact implementations for any shared procedures used by baselines.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes/No/Not Applicable]
      We specify our metrics in Section˜5 and state that we report results over $10$ random seeds.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes/No/Not Applicable]
      Experiments were run locally on a laptop. Details are given at the start of Appendix˜G.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes/No/Not Applicable]
      We declare dependencies on two previous repositories in Appendix˜G and cite the original works.
    2. (b)
      The license information of the assets, if applicable. [Yes/No/Not Applicable]
      We report the license in Appendix˜G and also in our source code.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes/No/Not Applicable]
      We provide the full code repository as supplemental material.
    4. (d)
      Information about consent from data providers/curators. [Yes/No/Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Yes/No/Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Yes/No/Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Yes/No/Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Yes/No/Not Applicable]
@@ A.1 The Distributional Definition of the Fokker-Planck Equation
In this setting, the existence and uniqueness of a weak solution follows from [Bogachev et al., 2022, Theorem 9.4.8, Example 9.4.7], which applies because $\sigma^{2}$ is constant and we assumed that $\nabla\Psi$ is Lipschitz and obeys the growth condition, $\|\nabla\Psi(x)\|\leq K(1+\|x\|)$. This result is also given in Stroock [2008, Theorem 1.1.9], which additionally shows that the family $(\mu_{t})_{t\in(0,T)}$ is continuous under the topology of weak convergence. Thus, the evolution of marginals can be identified with a continuous transition probability function, even under the weak formulation.
@@ A.2 Additional Proofs
The following results are used to complete the proof of ˜3.3. Since the observed marginals in ˜3.3 are sampled after some time $T_{1}>0$, we note that it suffices to consider times $t\geq T_{1}>0$. Also recall that for this result, we assume that $\Psi\in C^{\infty}(\mathbb{R}^{d})$. This facilitates the proof, by ensuring that any marginal $p(\cdot,t)$ of a candidate gradient-flow SDE exhibits nice regularity and decay. Indeed, by elliptic regularity theory, the Fokker-Planck operator $\mathcal{L}_{\shortminus\nabla\Psi,\sigma^{2}}$ smooths marginals within any finite time.
@@
As defined in ˜3.3, let $\mathcal{I}_{i}=\{t\in[T_{i},T_{i+1}]\mid p(\cdot,t)=q(\cdot,t)\}$, and let $\{t_{n}\}_{n\geq 0}$ be a sequence of times in $\mathcal{I}_{i}$, which converges to $t_{i}^{*}$. Then, $p(\cdot,t_{i}^{*})=q(\cdot,t_{i}^{*})$ and $\frac{\partial}{\partial_{t}}p(\cdot,t_{i}^{*})=\frac{\partial}{\partial_{t}}q(\cdot,t_{i}^{*})$.
@@
Let $(p(\cdot,t))_{t\geq T_{1}}$ be the marginals of a Langevin SDE (1) from the earliest possible observation time in the setting of ˜3.3, and suppose that $\Psi\in C^{\infty}(\mathbb{R}^{d})$. Then, the free energy $J(\rho)=\int_{\mathbb{R}^{d}}\Psi(x)\rho(x)\mathrm{d}x+\beta\int_{\mathbb{R}^{d}}\rho(x)\log\rho(x)\mathrm{d}x$, such that $\beta=\frac{\sigma^{2}}{2}$, strictly decreases in time, unless the process is at its stationary Gibbs distribution, $p_{\mathrm{eq}}$. Thus, $p_{\mathrm{eq}}$ is the only possible marginal that can repeat at distinct times.
@@
$\frac{\text{d}}{\text{d}t}J(p(\cdot,t))$ $=\int_{\mathbb{R}^{d}}\bigl(\Psi(x)+\beta(\log p(x,t)+1)\bigr)\,\frac{\partial}{\partial t}p(x,t)\,\mathrm{d}x$ (22)
@@ Appendix B IDENTIFIABILITY IN THE TIME-INHOMOGENEOUS CASE
We apply Theorem˜3.2 on each regime $t_{i},t_{i+1})$. Since the regimes form a partition of $[0,T]$, this guarantees that there is a unique SDE with the observed marginals $\bigl(p(\cdot,t\bigr)_{t\in[0,T]}$. ∎
@@ Appendix C ADDITIONAL NON-IDENTIFIABILITY EXAMPLES
For gradient-flow SDEs, we have proven that non-identifiability from marginals arises if and only if the observed marginals also correspond to the marginals of a residual process in equilibrium. An example for the time-homogeneous case was given in ˜2.2, such that the equilibrium marginals are constant, and an example for the time-inhomogenous case was given in ˜B.3, such that the equilibrium distribution of the residual process continuously changes in time.
@@
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}-10&0\\ 0&b\end{bmatrix}\mathrm{d}t+\begin{bmatrix}\sqrt{10}&0\\ 0&c\end{bmatrix}\mathrm{d}W_{t}.$ $Y_{0}\sim\mathcal{N}(0,1)\times[a,b]$ (30)
@@
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}Y_{t}\>\mathrm{d}t+\mathrm{d}W_{t}.$ $Y_{0}\sim\mathcal{N}(0,I)$
@@
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}1/3&4/3\\ 2/3&-1/3\end{bmatrix}Y_{t}\>\mathrm{d}t+\begin{bmatrix}1&2\\ -1&-2\end{bmatrix}\mathrm{d}W_{t}.$ $Y_{0}=\begin{bmatrix}1\\ -1\end{bmatrix}$
Appendix D ADDITIONAL NOTES ON NN-APPEX
D.1 Convergence
@@
By Girsanov’s theorem, (31) is only finite for $Q$ and $P$, if their underlying SDEs share the same diffusion [Vargas et al., 2021]. Besides non-identifiability, this is another cited reason for why previous Schrödinger Bridge methods assume that diffusivity must be fixed [Shen et al., 2025]. However, in the practical setting, we only observe a finite number of marginals, so we should instead optimize $\mathrm{KL}$ divergence over the discretized path space $\Omega_{N}=C(\{t_{i}\}_{i=0}^{N-1},\mathbb{R}^{d})$. Then, for each law on paths, we only observe the couplings, $Q_{t_{i},t_{i+1}}$ and $P_{t_{i},t_{i+1}}$, between consecutive times. Let $Q^{N}$ and $P^{N}$ denote the concatenations of these couplings from $t_{0}$ to $t_{N-1}$. Then, by Benamou et al. [2019][Lemma 3.4], evaluating the $\mathrm{KL}$ divergence over $\Omega_{N}=C(\{t_{i}\}_{i=0}^{N-1},\mathbb{R}^{d})$ yields
@@
As described in Section˜4, nn-APPEX will alternate between finding a law on paths $P$ that minimizes KL in the first argument (trajectory inference) and finding a law on paths $Q$ that minimizes KL in the second argument (MLE parameter estimation). Note that each optimization is subject to distinct hard constraints, which respectively stipulate that $P\in\Pi(p(\cdot,t_{i})_{i=0}^{N-1})$ adheres to the produced marginals and that $Q$ is the law of a gradient-flow SDE. The algorithm therefore performs alternating projections, onto distinct spaces. In fact, by ˜3.3, only one gradient-flow SDE $Q_{\shortminus\nabla\Psi,\sigma^{2}}$ obeys the marginal constraints, which implies that the intersection of the two projection spaces is a singleton. In other words,
$\inf\limits_{\begin{subarray}{c}P\in\Pi\bigl(p(\cdot,t_{i})_{i=0}^{N-1}\bigr)\end{subarray}}\mathrm{KL}(P\|Q_{\shortminus\nabla\Psi,\sigma^{2}})=0\iff P=Q_{\shortminus\nabla\Psi,\sigma^{2}}.$
By construction, the iterates will produce a monotonically decreasing sequence of $\mathrm{KL}$ divergences between reconstructed laws and MLE estimated laws. With the above argument, we have that the $\mathrm{KL}$ divergences are bounded above by some $M>0$, and from below by 0 (attained by the true gradient-flow SDE parameters $(-\nabla\Psi,\sigma^{2})$). Thus, the sequence of $\mathrm{KL}$ divergences must converge. However, as noted in Shen et al. [2025], this does not imply that the arguments necessarily converge, since multiple pairs $(P,Q)$ may produce the same divergence, such that the iterative algorithm gets stuck in a cycle. While this has not been observed empirically, and is impossible if the divergence is 0 (since there is a unique gradient-flow SDE satisfying the marginal constraints), it is an open question whether the iterative scheme will converge to the unique optimal parameters. Indeed, while convergence of the iterative SB refinement algorithm has been proven in the case where the family of admissible probability transition densities is convex [Shen et al., 2025, Proposition 1], this condition does not generally hold, and is false for the family of gradient-flow SDEs.
D.2 Runtime
@@
We note that nn-APPEX’s runtime will vary for different implementations of these subprocedures. In particular, there are multiple approaches for solving the multi-marginal SB problem. One approach is to consider all contiguous pairs of marginals, and then to apply iterative proportional fitting (Sinkhorn’s algorithm) on each pair to determine the distribution over couplings [Shen et al., 2025, Guan et al., 2024]. To enforce consistency across transitions, we instead apply a multi-marginal iterative proportional fitting across all marginals, see [Marino and Gerolin, 2020][(4.9)]. This algorithm is analogous to the two-marginals setting, but it instead rescales per-time slices for each marginal, rather than just the endpoints. We observed better accuracy for this approach, given the same stopping criteria and maximum number of iterations, and were thus able to reduce runtime with relatively fewer iterations. We similarly note that MLE drift estimation can be implemented using different neural network architectures and hyperparameters. In all experiments we use $128$ neurons per layer, but this can be changed with some alterations to both the flexibility and the runtime (Table 2).
Table 2: Average runtime of drift estimation step (14) using different neural network widths.
NN Width Avg Time (s)
16 1.32
32 1.85
64 3.20
128 6.57
256 13.20
In contrast to the drift estimation, the diffusion MLE estimate is a closed formula, and is thus cheap to compute. We report the average runtime (in seconds) of each of the three subprocedures for a single iteration, run on our main experiments in Table 3.
@@ Appendix E MAXIMUM LIKELIHOOD ESTIMATION FOR DRIFT AND DIFFUSION
If we observe $M$ independent trajectories over N1N-1 time steps ($N$ observed times), let $X=\bigl\{x_{i\Delta t}^{(m)},x_{(i+1)\Delta t}^{(m)}:m\in\{1,\cdots,M\},i\in\{0,\cdots,N-2\}\bigr\}$. We similarly observe that the negative log-likelihood function is given by
$l\left(\theta,\sigma^{2}|X\right)=-\log\left(\prod_{m=1}^{M}\prod_{i=0}^{N-2}p(x_{(i+1)\Delta t}^{(m)}|x_{i\Delta t}^{(m)})\right)=-\sum_{m=1}^{M}\sum_{i=0}^{N-2}\log(p(x_{(i+1)\Delta t}^{(m)}|x_{i\Delta t}^{(m)})),$
@@
$\ell(\theta)=\sum_{m=1}^{M}\sum_{i=0}^{N-2}\Big\|x_{(i+1)\Delta t}^{(m)}-x_{i\Delta t}^{(m)}+\nabla\Psi_{\theta}(x_{i\Delta t}^{(m)})\Delta t\Big\|_{2}^{2}.$ (35)
@@
$0$ $=\sum_{m=1}^{M}\sum_{i=0}^{N-2}\frac{d}{2\sigma^{2}}-\frac{1}{2\sigma^{4}\Delta t}\sum_{m=1}^{M}\sum_{i=0}^{N-2}\Big\|x_{(i+1)\Delta t}^{(m)}-x_{i\Delta t}^{(m)}+\nabla\Psi_{\hat{\theta}}(x_{i\Delta t}^{(m)})\Delta t\Big\|_{2}^{2}$